OpenStudy (anonymous):

HELP WILL GIVE MEDALS!! Write an explicit definition for the sequence -3,1,5,9,...

OpenStudy (anonymous):

@Hero can you help

OpenStudy (anonymous):

@mathmale can you help

myininaya (myininaya):

Hint: I wonder if this an arithmetic sequence.

OpenStudy (mathmale):

Hello, dmb97, Take a good look at the first few terms of this sequence: -3,1,5,9,... How do we obtain that 1 from -3? that 5 from 1?

OpenStudy (mathmale):

There are two primary types of sequences that we learn at this level: arithmetic and geometric. Which do you think is the case in THIS particular sequence?

OpenStudy (anonymous):

a arithmetic and its +4

OpenStudy (anonymous):

x = 4*n + 1

OpenStudy (mathmale):

So you'd start with -3, add 4 and get 1. Then you'd add 4 again and get 5. And so on. Looks like progress to me.

OpenStudy (mathmale):

Note that this is an ARITH. sequence. When n=1, the first term of the seq. is -1. When n=2 the second term of the sequence is 3. Can you write an arithmetic sequence that produces these results?

OpenStudy (mathmale):

\[a _{n}=???\]

OpenStudy (mathmale):

\[a _{1}=-1\]

OpenStudy (mathmale):

\[a _{2}=a _{1}+?\]

OpenStudy (mathmale):

Your formula must involve n and must involve 4 in some manner.

OpenStudy (mathmale):

Sorry, but I need to get off the Internet soon. Happy to help you, but need fast answers from you to stay engaged.

OpenStudy (anonymous):

a2=a1+4n

OpenStudy (mathmale):

or, more generally,\[a _{n+1{}}=a _{n}+4\]

OpenStudy (mathmale):

Check this out and determine whether or not it "works" for n = 1, 2, 3, 4, ... Unfortunately I have to get off the 'Net now. Hope this discussion has been of some help to you.

OpenStudy (mathmale):

Please let me know how this problem ended for you. Are you clear regarding what you were supposed to do, and with the formula we eventually obtained? This is called a "recursive" formula, because each term of the sequence depends upon the previous term. There's another way to write a formula for this arithmetic sequence.